OpenStudy (anonymous):

How many 2-inch cubes can be placed inside a larger cube that measures 2 feet on each side? a) 288 b)576 c)768 d)1728

OpenStudy (anonymous):

2 ft=2*12=24 in volume of bigger cube=24*24*24 in^3 volume of smaller cube=2*2*2 in^3 \[no.~ of~ cubes =\frac{ 24*24*24 }{ 2*2*2 }=?\]

OpenStudy (anonymous):

13824/8 = 1728 Thank you!

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

yw